# FLAF: Focal Line and Feature-constrained Active View Planning for Visual Teach and Repeat

### v2

@@ Abstract
This paper presents FLAF, a focal line and feature-constrained active view planning method for tracking failure avoidance in feature-based visual navigation of mobile robots. Our FLAF-based visual navigation is built upon a feature-based visual teach and repeat (VT&R) framework, which supports many robotic applications by teaching a robot to navigate on various paths that cover a significant portion of daily autonomous navigation requirements. However, tracking failure in feature-based visual simultaneous localization and mapping (VSLAM) caused by textureless regions in human-made environments is still limiting VT&R to be adopted in the real world. To address this problem, the proposed view planner is integrated into a feature-based visual SLAM system to build up an active VT&R system that avoids tracking failure. In our system, a pan-tilt unit (PTU)-based active camera is mounted on the mobile robot. Using FLAF, the active camera-based VSLAM operates during the teaching phase to construct a complete path map and in the repeat phase to maintain stable localization. FLAF orients the robot toward more map points to avoid mapping failures during path learning and toward more feature-identifiable map points beneficial for localization while following the learned trajectory. Experiments in real scenarios demonstrate that FLAF outperforms the methods that do not consider feature-identifiability, and our active VT&R system performs well in complex environments by effectively dealing with low-texture regions.
@@
VT&R, Active View Planning, Low Texture
@@ I Introduction
As a type of natural visual sensor, monocular cameras are cost-effective, energy-efficient, and versatile, making them suitable for a wide range of environments and applications. The T&R approaches that predominantly utilize visual sensors are referred to as Visual Teach and Repeat (VT&R)[1], which is a significant motivation of the research in visual simultaneous localization and mapping (VSLAM)[5]. Our active VT&R system integrates a pan-tilt unit (PTU)-based active camera with feature-based VSLAM (see Fig. 1). During the teaching phase, the mobile robot reconstructs the surrounding landmarks while traversing a path under guidance[6]. In the repeat phase, the previously saved path map is reloaded to localize the robot for navigating the taught trajectory.
@@
Figure 1: The left side of Fig. (a) illustrates the failure of the feature-based VT&R system to learn a complete path due to a low-texture region. The corresponding orientation and view of the passive camera are depicted in Fig. (b). In contrast, with our FLAF method and an active camera, the VT&R system successfully navigates this challenging area by directing the camera toward feature-rich regions, as shown in Fig. (c). Fig. (d) presents our mobile robot equipped with an active camera.
@@
Although feature-based VSLAM systems[7, 8] achieve impressive robustness and stability in large indoor and outdoor environments[9], they all suffer from the view angle-dependent affine change[10] of features, and the map for real-time localization are usually too sparse for other complex applications[11]. Fortunately, the limited view-angle invariance of features and sparse map for reliable localization is enough for path following in VT&R. In this work, we achieved a high success rate and reliability in passive VT&R (using a fixed camera) with feature-based monocular VSLAM. However, textureless regions in human-made environments are still limiting VSLAM and VT&R to be used in the real world. To solve this problem, existing active SLAM works mostly choose to change the robot trajectory, which is not suitable for VT&R[12]. In order to use an active camera to actively select informative views without interfering in the T&R trajectory, our work focuses on designing an active view planning method with active camera-based VSLAM suitable for VT&R.
@@
Our motivation is to design a feature-based VT&R system coupled with active view planning for tracking failure avoidance. Previous active camera-based VSLAM systems[14, 15, 16] have primarily focused on maintaining stable localization during the mapping process but have not demonstrated how to effectively reuse the map with active view planning for navigation tasks. The key difference between mapping and reusing the map is that there are more choices of map points when the map is complete during navigation, compared to the mapping process when the active camera has no choice but to orient to the newly built map points. As shown in Fig. 2, existing methods[14, 15, 16] that rotate the active camera to fixate on more map points that are beneficial for tracking may fail in some VT&R cases because they do not consider the feature identifiability[25], which refers to the ability of the feature detector to identify a 3D map point. While the robot is moving forward in the direction indicated by the arrows (see Fig. 2), existing view planning methods mostly focus the active camera on the regions dense with map points. In many situations, these map points targeted by the active camera are triangulated by earlier keyframes taken from viewpoints significantly different from the current one, making them unidentifiable by the feature algorithms due to their substantially different visual appearances. To address this, we have designed an active camera-based VT&R system featuring an innovative active view planning method that accounts for the view angles of map points. Our FLAF-based active view planning is focal line-centric, as its direction dictates the angles of the active camera.
In this paper we present a visual navigation system with the PTU and a novel active view planning method for feature-based VT&R. The contributions of our work are summarized as follows: (1) We combine VSLAM, motion control in repeat, and feature-based active view planning to overcome the low-texture problem. To the best of our knowledge, this is the first demonstration of VT&R system coupled with active view planning. (2) We propose an online focal line and feature (FLAF)-constrained active view planning method that rewards the camera orientations that observe more feature identifiable map points at their mean-view directions and is suitable for both phases of VT&R.
@@ II RELATED WORK
Our active VT&R is based on VSLAM system tightly coupled with active view planning for tracking failure avoidance.
@@ II-A Visual Teach and Repeat
The classical work of VT&R[1] using a feature-based stereo VSLAM was later developed as VT&R2[2], which utilizes multiple taught experiences to address environmental appearance change. Despite the significant progress, this series of works all suffer from the tracking failure caused by low-texture regions[12] as they all rely on fixed cameras and feature-based VSLAM. Based on the well-established VT&R2[2], Warren et al.[17] build a gimbal-stabilized VT&R system in which the gimbal is passively utilized to stabilize the camera or manually steered to avoid degeneracy in the teach phase. In the repeat phase, the camera actively rotates to the nearest keyframe in the taught graph. Although this work justified the necessity of an active camera in VT&R, the gimbal is manually steered in the teach phase instead of autonomous operation. Previous works designed various active view planning methods[14, 15, 16, 17] for active camera-based VSLAM. However, the recently proposed uncertainty-driven view planning (UDVP)[16] that rotates the camera toward more nearby map points easily fails in the repeat phase for it ignores the affine change of features[10] (Fig. 2). To solve these problems, we implement the same autonomous view planning method (FLAF) in both phases of VT&R.
Figure 2: This is a failure case of existing methods considering only the quantity and distance of map points to the camera center. The keyframes are selected by the VSLAM system in the teach phase. Although the blue points are denser and closer to the camera, they were triangulated from significantly different viewpoints in earlier keyframes, making them difficult to detect from the current viewpoint.
Faced with the aforementioned tracking failure and occlusion, Mattamala et al. [13] designed a VT&R system allowing the quadruped to switch between multiple cameras set on different positions of the robot. However, each one among the cameras faces the same problem of tracking failure. Every camera builds several sub-maps where completion and consistency are difficult to guarantee in sub-map merging without well dealing with the challenging views. Our proposed active VSLAM can help each one of the multiple cameras in[13] to improve the performance of its system. In the meanwhile, our VT&R system is more natural and compact to use fewer computing resources.
@@ II-B Visual Simultaneous Localization and Mapping
Our VT&R system consists of a 3D reconstruction module for the mobile robot to remember a traversed path. In the teach phase, the robot learns the landmarks in feature-based VSLAM. With the input of an image set that shares observations of the environment, the task of estimating camera motions and a geometrical reconstruction is called Structure From Motion (SfM)[18]. For a mobile robot with a moving video camera, a system that performs SfM for every image as it is captured is called real-time SfM or VSLAM. Specifically, our proposed active view planning method is designed according to the local map built by VSLAM systems[9].
The most popular implementations of VSLAM[7, 9, 8] follow the principle of aligning the current image to the already-built map for camera localization. The data association between current frame and keyframes and bundle adjustment (BA) are conducted locally and globally to obtain a consistent map[19]. To use time-consuming bundle adjustment in SLAM for optimizing a consistent map, a hierarchical optimization strategy is proposed with the concept of window and local bundle adjustment[20].
In [9], a sophisticated local map is designed to align the current frame and implement the local BA. Motion estimation by aligning the current image to the local map is called local map tracking[7]. Deng[12] et al. identify a particular tracking failure caused by the incapability of associating enough features. In [12] the authors also indicate that the possibility of tracking failure approximates zero if the associated map points is more than a level.
@@ II-C Active View Planning for VSLAM
Seminal work of active view planning for VSLAM is [15], in which Davison et al. indicates the solution of active vision in navigation is using serial views on a succession of features for stable localization and deciding the time to explore new features. Following this idea, [14] proposes to observe the unexplored regions and reset the active camera if features are not enough for localization. However, this strategy doesn’t guarantee an accurate localization because the view direction same to the initial one doesn’t capture the same initial image as the robot is moving.
In active camera-based VSLAM, the primary problem before the exploration of new features is trying its best to maintain a stable and accurate localization[15]. In [16] the authors propose the UDVP method to measure the quality of the camera angles with respect to every map point. This UDVP model shows fine performance in capturing more existing map points in the Field of View (FoV) of the camera. However, it (shown in Fig. 2) doesn’t consider the feature identifiability in the repeat phase which results in tracking failure by looking at unrecognizable map points. Our FLAF method uses the observation model shown in Fig. 3, which utilizes the angle between the camera focal line and the line of sight to a map point to make the camera capture more built map points. In the meanwhile, the angle between the imaging light path of the point and the normal line of the map point is measured to qualify the feature identifiability.
@@
The idea to consider both the angles of $\alpha_{1}$ and $\alpha_{2}$ shown in Fig. 3 was first described in the VSLAM system of[9] for feature selection, which serves as the foundation for our system. This strategy is also utilized in the active SLAM method presented in [12], which employs a fixed camera for both navigation and exploration. The model in [12] defines specific ranges of distance and $\alpha_{2}$ to screen map points within the camera’s frustum. The “FLAF without scoring” method in our comparison experiments can be seen as an implementation of active VT&R using the model from [12], despite the differences in the robot and task. It is noteworthy that Mostegel et al.[25] identified several metrics for feature recognition and validated the effectiveness of using a cosine function to measure the feature recognition probability, which indicates the likelihood of feature identification from different view angles.
Figure 4: This is the framework of our active camera-based VSLAM system, which serves as the foundation for the active VT&R system. Our active view planning module is integrated between the local map tracking module and the local mapping module in the passive VSLAM system. The PTU angles in the range of 60 for both yaw angle and pitch angle are sampled with a set interval of 2 and evaluated by our FLAF-constrained view planner. The pan-tilt sample with the highest score, as determined by FLAF, is sent to the PTU control module to adjust the camera direction accordingly.
@@ III Approach
We build our VT&R system on the active camera-based VSLAM framework shown in Fig. 4, which integrates ORB-SLAM2[9] and our FLAF-constrained active view planning. Our method for active camera-based path following is similar to that described in [17], which resolves the robot poses from camera localization and PTU angles as the input of VT&R.
To increase the number of map points in the central area of the FoV, we designed a scoring function (Eq. (3)) based on $\cos({\alpha_{1}})$ (see Fig. 3) to rotate the camera toward regions dense with map points. As the UDVP model[16] uses a sigmoid function of $\alpha_{1}$, we apply a cosine function to reduce the computational cost. By adding up the scores of all map points within the FoV, each one of them contributes to the total score of this PTU sample. Map points with smaller $\alpha_{1}$ values receive higher scores, which encourages the clustering of map points in the central area of the FoV. Consequently, samples of the active camera that capture more map points with smaller $\alpha_{1}$ in their FoV achieve higher scores.
Another goal of our design is to score the map points with good feature identifiability relative to the samples of PTU angles. Feature identifiability refers to the ability of the feature detector to recognize the map point from a specific position and orientation. With the path map built, the VT&R system can determine the existence of a map point at a 3D coordinate. But if the robot rotates the camera toward this map point from an arbitrary position, there is a probability[25] that this map point may not be detected in the current image due to affine changes[10]. We assume the keyframes that observe the same map point define a view distribution within which this map point can be identified. The mean viewing direction is considered the normal of the map point, around which the view distribution is defined. This definition of the map point normal was also used in [9, 12]. Mostegel et al.[25] justified using $\cos{\alpha_{2}}$ as the metric of the probability of feature identification to account for the observation of a feature from different viewpoints and view angles. Our scoring function, $\cos({\alpha_{1}})\cdot\cos({\alpha_{2}})$, multiplies two cosine functions to prioritize map points that score highly on both metrics.
III-A VSLAM with Active Camera
In the classical feature-based VSLAM[9], on which we build our active camera-based SLAM system, the next frame is passively decided by the trajectory of the robot. To avoid low-texture views, we make passive VSLAM tightly coupled with a feature-based active view planning module. Specifically, we insert (Fig. 4) the active view planning module between the local map tracking and the local mapping as they are separated into different threads in implementation.
@@ III-A VSLAM with Active Camera
While the mobile robot is moving along a path, images with equally spaced timestamps are sequentially captured by the moving camera and sent to the VSLAM system. For every input image $\mathbf{I}_{k}:\mathbb{R}^{2}\rightarrow{\mathbb{R}}$, ORB features[21] are extracted and sent to the local-map tracking module, in which features are aligned with the local map to estimate current camera pose $\mathbf{X}_{k,w}\in\rm{SE(3)}$ in world coordinate. This process of computing a camera pose according to the local map is called local map tracking. If enough new features are found, this frame is decided as a keyframe and these features are triangulated into the map based on data association between local keyframes. This process of projecting new keyframes and features into the map space (world coordinate) is called local mapping.
A local map including a network of keframes $\big{\{}\boldsymbol{F}_{i}\big{\}}$ connected according to feature matching and thier associated map points $\big{\{}\mathbf{P}_{j}^{i}|\mathbf{P}_{j}^{i}\in\mathbb{R}^{3},j=0,1,2,...,m_{i}\big{\}}$ are denoted with:
@@
where $m_{i}$ is the quantity of the map points associated with the i-th keyframe and $n_{i}$ is the quantity of the keyframes in the local map. A map point only refers to a 3D coordinate in map space, but its descriptor can be computed from the keyframe by which the point is triangulated. According to the descriptor matching, keyframes share observations with $\mathbf{I}$ to make up the local map with their associated map points.
For any $\mathbf{P}_{j}^{i}$ in $\boldsymbol{M}_{l}$ that is within the visible frustum of the current camera pose, if it can be matched with an ORB feature at $\mathbf{p}_{j}^{i}\in\mathbb{R}^{2}$ on $\mathbf{I}_{k}$, it is reprojected onto the $\mathbf{I}_{k}$ by pinhole camera model $\pi:\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$ and get the reprojected coordinate $\pi(\mathbf{P}_{j}^{i})\in\mathbb{R}^{2}$. The reprojection error as in Eq. (2) is applied for optimizations in VSLAM:
@@
After local map tracking, we have the local map and current camera pose needed in active view planning. Based on the local map and current camera pose, the next best angles of the PTU can be decided. Thus we achieve using current partial perception results as feedback to control the sensor and image input. According to the results in [12], the pose estimation shown by Equation (3) fails if the magnitude of $S$ is below a threshold. Therefore, [16] proposes a view planner to increase $N_{S}$. To achieve this by our sampling-based optimization, the pan-tilt angles are sampled as $\mathbf{q}=(pan,tilt)$ and transformed into $\mathbf{T}_{pt}(\mathbf{q})\in\rm{SE}(3)$ to obtain the corresponding sample of camera pose $\mathbf{X}_{k}^{\prime}$, which is directly evaluated:
$\mathbf{X}_{k}^{\prime}=\mathbf{T}_{pt}(\mathbf{q})\ast\mathbf{X}_{k}$ (4)
while the robot is moving along a path, the PTU samples are evaluated by view planners at a set frequency, and the best one is sent to the PTU to rotate the active camera. During the simultaneous movements of the PTU and the robot, the images input to the VSLAM are decided by the PTU angles and the robot trajectory, while the PTU angles are decided by the robot trajectory and the local map of the VSLAM system.
III-B Active View Planning for Feature-based VT&R
We use Eq. (5) to represent the magnitude of map points within the visible frustum of the camera that can be matched with an ORB feature in $\mathbf{I}_{k}$:
$N_{S}=f\left(\mathbf{X}_{k},\boldsymbol{M}_{l}\right)$ (5)
Eq. (5) indicates that $N_{S}$ is a function of the current camera pose and the local map. Using the UDVP observation model[16], we control the active camera to see more and closer map points. Given a PTU sample, the UDVP[16] model scores map points with a smaller distance to the camera center and a smaller view angle between the line of sight to them and the camera’s focal line. Thus $N_{S}=f(\mathbf{X}_{w})$ is increased and tracking failure probability is decreased compared to passive VSLAM. This effect can be represented by the inequality:
$f(\mathbf{T}_{pt}\mathbf{X}_{k})>f(\mathbf{X}_{k})$ (6)
However, in the repeat phase, a complete map of the taught path is built. Although there exist more local map points that are close to the optical center and have a small angle between $OP$ and $n_{c}$ ($\alpha_{1}$ shown in Fig. 3), some of them can not be identified by the feature detector due to a big angle between the focal line and mean view angle of the map point. This phenomenon can be summarized as “looking at points visible but not identifiable”, which results in a higher failure rate of repeat. To address this, we apply the observation model shown in Fig. 3. One distance and two angles are measured for every PTU sample and map point to support our view planner:
  • Scoring the pan-tilt sample that makes the map point in the range of $(d_{1},d_{2})$ which is the scale invariance range of the image pyramid.
  • Scoring smaller $\alpha_{1}(\mathbf{T}_{pt}\mathbf{X}_{k},\mathbf{P}_{j}^{i})$ shown in Fig. 3 which refers to the angle between camera’s focal line $n_{c}$ and ray line $OP$ of the point.
  • Scoring reward smaller $\alpha_{2}(\mathbf{X}_{k},\mathbf{P}_{j}^{i})$ shown in Fig. 3 which refers to the angle between mean view line $n_{p}$ and ray line $OP$ of the point.

### v3

@@ Abstract
This paper presents FLAF, a focal line and feature-constrained active view planning method for autonomous orientation adjustment of a rotatable active camera during mobile robot navigation. FLAF is built on a visual teach-and-repeat (VT&R) system, which enables robots to cruise various paths that fulfill many daily autonomous navigation requirements. The VT&R system integrates Visual Simultaneous Localization and Mapping (VSLAM) with trajectory following. However, tracking failures in feature-based VSLAM, particularly in textureless regions common in human-made environments, poses a significant challenge to real-world VT&R deployment. To address this, the proposed view planner is integrated into a feature-based VSLAM system, creating an active VT&R solution that mitigates tracking failures. Our system features a Pan-Tilt Unit (PTU)-based active mounted on a mobile robot. Using FLAF, the active camera-based VSLAM (AC-SLAM) operates during the teaching phase to construct a complete path map and in the repeating phase to maintain stable localization. FLAF actively directs the camera toward more map points to avoid mapping failures during path learning and toward more feature-identifiable map points while following the learned trajectory. Experimental results in real scenarios show that FLAF significantly outperforms existing methods by accounting for feature identifiability, particularly the view angle of the features. While effectively dealing with low-texture regions in active view planning, considering feature identifiability enables our active VT&R system to perform well in challenging environments.
@@
VT&R, Active View Planning, Visual SLAM
@@ I Introduction
As a type of natural visual sensor, monocular cameras are cost-effective, energy-efficient, and versatile, making them suitable for a wide range of environments and applications. The T&R approaches that predominantly utilize visual sensors are referred to as visual teach-and-repeat (VT&R)[1], which is a significant motivation of the research in visual simultaneous localization and mapping (VSLAM)[5]. During teaching, the robot reconstructs the surrounding landmarks by VSLAM while traversing a path under guidance[6]. In the repeating phase, the previously saved path map is reloaded to localize the robot for navigating the taught trajectory.
@@
Figure 1: (a) shows the failure of the passive VT&R system to learn a complete path due to a low-texture region, which is overcome by our active VT&R with view planning. The orientations and views of the fixed camera are depicted in (b). In contrast, with our FLAF-constrained view planning and an active camera, the active VT&R system successfully navigates this challenging path by directing the camera toward feature-rich regions, as shown in (c). (d) presents our mobile robot equipped with a PTU-based active camera.
@@
Although feature-based VSLAM systems[7, 8] achieve impressive robustness and stability in large indoor and outdoor environments[9], they all suffer from the view angle-dependent affine changes[10] of features, and the map for real-time localization are usually too sparse for complex applications[11] that need dense reconstruction. Fortunately, the limited view-angle invariance of features and the sparse map for reliable localization are enough for path following in VT&R. Using a passive VSLAM system, we achieved a high success rate and reliability in passive VT&R (using a fixed camera) with feature-based monocular VSLAM. However, tracking failures[12] caused by textureless regions in human-made environments are still limiting VSLAM and VT&R to be used in the real world. To solve this, existing active SLAM methods[12, 25] mostly choose to change the robot trajectory, which is not suitable for VT&R. To actively select informative views without interfering in the trajectory, our active VT&R system integrates a pan-tilt unit (PTU)-based active camera with feature-based VSLAM (Fig. 1).
@@
Our motivation is to design a feature-based VT&R system coupled with active view planning for tracking failure avoidance. Previous active camera-based VSLAM (AC-SLAM) methods[14, 15, 16] have primarily focused on maintaining stable localization during the mapping process but have not demonstrated how to effectively reuse the map with active view planning for navigation tasks. Compared to the mapping process when the active camera has no choice but to orient to the newly built map points, there are more choices of map points when the map is complete during navigation. While the robot is moving forward in the direction indicated by the arrows (Fig. 2), existing view planning methods [14, 15, 16] that focus the active camera on the regions dense with map points may fail in some VT&R cases because they do not consider the feature identifiability (the ability of the feature detector to identify a 3D map point). These map points targeted by the active camera may be triangulated by earlier keyframes taken from viewpoints significantly different from the current one, making them unidentifiable by the feature algorithms due to their substantially different visual appearances[25]. To address this, we have designed an active camera-based VT&R system featuring an innovative active view planning method that accounts for the view angles of map points. Our FLAF-based view planner is focal line-centric, as its direction dictates the angles of the active camera.
In this paper, we present a feature-based active VT&R system incorporating an active camera and a novel active view planning method. The main contributions are as follows: (1)We propose a focal line and feature (FLAF)-constrained view planner that addresses failures of visual repeating by accounting for the angle difference between the current viewpoint and those at which the map points were triangulated. (2) We integrate active view planning into passive VT&R to build up the active VT&R by resolving the robot poses from camera localization and PTU angles as the input of passive VT&R. To the best of our knowledge, this is the first demonstration of the VT&R system coupled with active view planning. (3) Our code is publicly available at: https://github.com/Changfei-Fu.
@@ II RELATED WORK
Our active VT&R builds on the VSLAM system tightly coupled with active view planning for tracking failure avoidance.
@@ II-A Visual Teach and Repeat
The classical work of VT&R[1] using a feature-based stereo VSLAM was later developed as VT&R2[2], which utilizes multiple taught experiences to address environmental appearance changes. Despite the significant progress, this series of works all suffer from the tracking failure caused by low-texture regions[12] as they all rely on fixed cameras and feature-based VSLAM. Based on the well-established VT&R2[2], Warren et al.[17] build a gimbal-stabilized VT&R system in which the gimbal is passively utilized to stabilize the camera or manually steered to avoid degeneracy in the teaching phase. In the repeating phase, the camera actively rotates to the nearest keyframe in the taught graph. Although this work justified the necessity of an active camera in VT&R, the gimbal is manually steered in the teaching phase instead of autonomous operation. Previous works designed various active view planning methods for AC-SLAM[14, 15, 16, 17]. However, they fail to account for the appearance change of the same feature observed from different view angles. Similar to the recently proposed uncertainty-driven view planning (UDVP)[16], these approaches tend to rotate the camera toward nearby map points, leading to frequent failures during repeating for they ignore the affine change of features[10] (Fig. 2). To solve these problems, we implement the same autonomous view planning method (FLAF) in both phases of VT&R.
Figure 2: The failure case of existing view planning methods that consider only the number of map points within the Field of View and their distance to the camera center. The keyframes are selected by the VSLAM system during teaching. During repeating, although the blue points are denser and closer to the active camera, they were triangulated from significantly different viewpoints of earlier keyframes, making them difficult to identify from the current view due to appearance change.
Confronted with the aforementioned tracking failure and occlusion, Mattamala et al. [13] designed a VT&R system that allows the quadruped to alternate between multiple cameras mounted at different positions on the robot. However, each one of the cameras still encounters the issue of tracking failure. Each camera builds several sub-maps, and ensuring their completion and consistency during merging is difficult without effectively addressing the challenging views. Our proposed AC-SLAM VSLAM can enhance the performance of each camera in this system[13]. Additionally, our VT&R system is designed more intuitive and compact, utilizing fewer computational resources.
@@ II-B Visual Simultaneous Localization and Mapping
Our VT&R system consists of a 3D reconstruction module for the mobile robot to remember a traversed path. In the teaching phase, the robot learns the landmarks in feature-based VSLAM. With the input of an image set that shares observations of the environment, the task of estimating camera motions and a geometrical reconstruction is called Structure From Motion (SfM)[18]. For a mobile robot with a moving video camera, a system that performs SfM for every image as it is captured is called real-time SfM or VSLAM. Specifically, our proposed active view planning method is designed according to the local map built by VSLAM systems[9].
The most popular implementations of VSLAM[7, 9, 8] follow the principle of aligning the current image to the already-built map for camera localization. Bundle adjustment (BA) based on data association between the current frame and keyframes is conducted locally and globally to obtain a consistent map[19]. To use time-consuming bundle adjustment in SLAM for optimizing a consistent map, a hierarchical optimization strategy is proposed with the concept of window and local bundle adjustment[20].
In [9], a sophisticated local map is designed to align the current frame and implement the local BA. Motion estimation by aligning the current image to the local map is called local map tracking[7]. Deng[12] et al. identify a particular tracking failure caused by the incapability of associating enough features. In [12], the authors also indicate that the likelihood of tracking failure approximates zero if the number of associated map points exceeds a certain threshold.
@@ II-C Active View Planning for VSLAM
The seminal work in active view planning for VSLAM is presented in [15], where Davison et al. demonstrate that the key to active vision for VSLAM is the continual views on a succession of features, along with determining the moment to explore new features. Following this concept, [14] suggests observing the unexplored areas and repositioning the active camera when there is an insufficient number of features for localization. However, this approach does not ensure precise localization, as a view direction identical to the initial one may not reproduce the original image due to the robot’s movement.
In active camera-based VSLAM, the primary challenge before exploring new features is maintaining stable and accurate localization[15]. In [16], Warren et al. propose the UDVP view planner to assess the pan-tilt angles relative to map points. The UDVP view planner performs well in capturing more map points within the camera’s Field of View (FoV). However, as shown in Fig. 2, it leads to tracking failure by orienting at unrecognizable map points during VT&R. Our FLAF view planner improves upon previous methods by the observation model shown in Fig. 3, which accounts for not only the view angle between the camera’s focal line and the light path of a map point but also the angle between the light path and the normal of map point, ensuring both the number of map points and feature identification.
The strategy to consider both $\alpha_{1}$ and $\alpha_{2}$ shown in Fig. 3 was initially introduced for feature selection in the VSLAM method described in [9], which forms the foundation of our active VT&R. This strategy is also employed in the active SLAM method presented in [12], where a fixed camera is used for navigation and exploration. Similar to [9], the method in [12] defines specific ranges of distance and $\alpha_{2}$ (see Fig. 3) to screen map points within the camera’s FoV. The “FLAF without scoring” method in our comparative experiments can be seen as an implementation of active VT&R using the view planner in [12], despite the differences in the robot and task. It is worth noting that Mostegel et al.[25] identified several metrics for feature recognition and validated the effectiveness of using a cosine function to estimate the likelihood of feature identification from different view angles.
@@ III Approach
III-A System Overview
Figure 4: Active VT&R pipeline, showing the relationship between VSLAM, view planning, and active VT&R. In both phases of VT&R, the AC-SLAM is performed and the PTU rotates automatically based on real-time perception, while the mapping module is deactivated during repeating.
As shown in Fig. 4, our active VT&R system is built on the AC-SLAM, which integrates ORB-SLAM2[9] and our FLAF-constrained active view planning. Our method for active camera-based path following resolves the robot poses from camera localization and pan-tilt angles, as the input of VT&R. During both phases of VT&R, the AC-SLAM operates continuously, with the camera orientation automatically adjusted in real-time based on perception feedback. The path map and robot trajectory created during teaching are reloaded before repeating, with the mapping module deactivated. The robot is initially placed near the taught path with a similar orientation to the taught one. Once the repeating begins, the robot autonomously cruises along the taught path. During repeating, the AC-SLAM provides the path-following module with trajectory reference and localization service, using the previously stored map.
III-B VSLAM with Active Camera
In passive VSLAM, the next image input is determined by the camera movement. To avoid low-texture views, we insert (Fig. 5) the feature-based active view planning between local map tracking and local mapping. During both phases of T&R, the AC-SLAM is performed, and the active camera rotates automatically based on real-time perception. The camera orientation, in turn, dictates the input images for the VSLAM.
@@ III-B VSLAM with Active Camera
While the camera is moving, images are sequentially captured and input to the AC-SLAM. For the k-th input image $\mathbf{I}_{k}:\mathbb{R}^{2}\rightarrow{\mathbb{R}}$, ORB features[21] are extracted for local-map tracking, where features are aligned with the local map to estimate current camera pose $\mathbf{X}_{k,w}\in\rm{SE(3)}$. If sufficient new features are found, this frame is decided as a keyframe and these features are triangulated into the map by local keyframes. This process of projecting new keyframes and features into the map coordinate is called local mapping. A local map including a network of keyframes $\big{\{}\boldsymbol{F}_{i}\big{\}}$ connected by feature matching and their associated map points $\big{\{}\mathbf{P}_{j}^{i}|\mathbf{P}_{j}^{i}\in\mathbb{R}^{3},j=0,1,2,...,m_{i}\big{\}}$ are denoted with:
@@
where $m_{i}$ is the quantity of the map points associated with the i-th keyframe and $n_{i}$ is the quantity of the keyframes in the local map. A map point only refers to a 3D coordinate, and its descriptor can be computed from the keyframe by which this point is triangulated. According to the descriptor matching, keyframes that share observations with $\mathbf{I_{k}}$ and their associated map points make up the local map.
For any $\mathbf{P}_{j}^{i}$ in $\boldsymbol{M}_{l}$ that is within the current FoV, if it can be matched with an ORB feature at coordinate $\mathbf{p}_{j}^{i}\in\mathbb{R}^{2}$ on $\mathbf{I}_{k}$, it is reprojected onto the $\mathbf{I}_{k}$ by pinhole camera model $\pi:\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$. The reprojection error as in Eq. (2) is applied for optimizations in VSLAM:
@@
After local map tracking, a sampling-based optimization (Fig. 5) of the next best view is adopted based on the local map and real-time localization. The pan-tilt angles are sampled as $\mathbf{q_{s}}=(pan,tilt)$ and transformed into $\mathbf{T}_{pt}(\mathbf{q_{s}})\in\rm{SE}(3)$ to obtain the corresponding sample of camera pose $\mathbf{X}_{S}$, which is directly scored by view planners at a set frequency:
$\mathbf{X}_{S}=\mathbf{T}_{pt}(\mathbf{q_{s}})\ast\mathbf{X}_{k}$ (4)
The best sample of pan-tilt angles is sent to the PTU to rotate the active camera. Thus we achieve rotating the active camera on real-time perception feedback.
Figure 5: The implementation of sampling-based view planning, which is inserted between the local map tracking and the local mapping of the passive VSLAM system. The PTU angles in the range of 60 for both yaw angle and pitch angle are sampled with a set interval of 2 and scored by our FLAF-constrained view planner. The pan-tilt sample with the highest score, as determined by FLAF, is sent to the PTU control module to adjust the camera direction accordingly.
III-C Active View Planning for Feature-based VT&R
To implement FLAF by sampling-based optimization, we adopt three metrics indicated in Fig. 3 for sample evaluation. For each pan-tilt sample and map point, one distance and two angles are measured to determine the next best view:
  • Reward the pan-tilt sample that places the map point within the feature-invariant distance range of $(d_{1},d_{2})$ relative to the camera.
  • Reward smaller $\alpha_{1}(\mathbf{X}_{S},\mathbf{P}_{j}^{i})$ shown in Fig. 3 which refers to the angle between camera’s focal line $n_{c}$ and light path $OP$ of the point.
  • Reward smaller $\alpha_{2}(\mathbf{X}_{k},\mathbf{P}_{j}^{i})$ shown in Fig. 3 which refers to the angle between mean view line $n_{p}$ and light path $OP$ of the point.